• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Meziani

    Log entry time 23:44:31 on February18,2011

    Entry number 347540

    keyword=Q2 and Q3 Right arm magnet report

    Q2 right arm triped sometimes during the day shit
    Noticed it while filling the shift checklist of the swing shift
    Could not be reset from the general tools in the counting house.
    Asked for access at 10:20 pm to reset it downstairs.
    Larry Selvy and Blaine Norum reset Q2 downstairs.
    Brought the Quad back to 1400 A first and then to its nominal value of 1269.5A

    Now it is up and running.

    Also noticed that the value of Q3 in the right arm was set to more than 1100 A. Before starting run 1531 it was set back to the same value as for the hydrogen runs of 02/17/2011 (1029.2 A).